OpenStudy (briellshaw):

If Rhymira had 16 books and 18 book marks in school when she came home she gave 7 books and 9 bookmarks away how many books and bookmarks would she have in all leftover?

OpenStudy (anonymous):

she would have 9 books and 9 bookmarks

OpenStudy (anonymous):

Take 16 books and subtract 7 from them. Take 18 book marks and subtract 9. Add those two answer you get together to get the total amount of books and bookmarks leftover.

OpenStudy (briellshaw):

18 IN ALL @Longjonsliver gave me the wrong answer this was a trick question you need to start paying attention more

OpenStudy (anonymous):

i gave you the right answer can you not add?

OpenStudy (anonymous):

he gave you the correct answer

OpenStudy (anonymous):

thankyou @randomkatey

OpenStudy (anonymous):

No problem @Longjonsliver

OpenStudy (briellshaw):

read the question again @randomkatey @Longjonsliver If Rhymira had 16 books and 18 book marks in school when she came home she gave 7 books and 9 bookmarks away how many books and bookmarks would she have in all leftover? how many books and bookmarks would she have in all IN ALL

OpenStudy (anonymous):

in total she has 9 of either which is 18 IN ALL

OpenStudy (anonymous):

16-7=9 18-9=9 9+9=18...

OpenStudy (anonymous):

If there were 7 books given away from 16 she would be left with 9 books. If there were 9 bookmarks given away from 18 book marks she would be left with 9 bookmarks. In total she has 18 books and bookmarks....to get that total you add the remaining books and book marks together....so if you add 9 books and 9 bookmarks together you get 18 books and bookmarks....make sense?

OpenStudy (anonymous):

@BriellShaw did you understand what I explained to you?

OpenStudy (anonymous):

Yes @longjonsliver you are not supposed to give the answer but help the person understand how to get the answer on their own. @BriellShaw see my explanation.

OpenStudy (anonymous):

i didnt give her the answer @OdinMW i can her a simple-er way to achieve the answer